# Towards Global AI Inclusivity: A Large-Scale Multilingual Terminology Dataset

### v1

@@ Abstract
The field of machine translation has achieved significant advancements, yet domain-specific terminology translation, particularly in AI, remains challenging. We introduced GIST, a large-scale multilingual AI terminology dataset containing 5K terms extracted from top AI conference papers spanning 2000 to 2023. The terms were translated into Arabic, Chinese, French, Japanese, and Russian using a hybrid framework that combines LLMs for extraction with human expertise for translation. The dataset’s quality was benchmarked against existing resources, demonstrating superior translation accuracy through crowdsourced evaluation. GIST was integrated into translation workflows using post-translation refinement methods that required no retraining, where LLM prompting consistently improved BLEU and COMET scores. A web demonstration on the ACL Anthology platform highlights its practical application, showcasing improved accessibility for non-English speakers. This work aims to address critical gaps in AI terminology resources and fosters global inclusivity and collaboration in AI research.
@@
A Large-Scale Multilingual Terminology Dataset
@@ 1 Introduction
Figure 2: Overview of GIST creation. AI terminologies were extracted from awarded papers using an LLM and combined with existing terminology dictionaries. After data cleansing, translations into five languages were generated via crowdsourcing, and an LLM was used to select the best candidate translation. Dataset quality was evaluated against two baselines: majority-vote-selected translations and translations from the 60-60 evaluation set, demonstrating the superior quality of GIST.
@@
The field of machine translation has made significant progress, with state-of-the-art models excelling across diverse tasks (Brown et al., 1990; Wu et al., 2016; Goyal et al., 2022; Haddow et al., 2022) and demonstrating effectiveness in translating between high-resource and low-resource languages (Yao and Wan, 2020; Costa-jussà et al., 2022; Ranathunga et al., 2023). Despite these successes, translating domain-specific scientific texts remains a persistent challenge, particularly for terminology translation (Cabré, 2010; Shuttleworth, 2014; Naveen and Trojovskỳ, 2024). General-purpose translation systems often falter in accurately translating terminologies, leading to loss of critical details (Dagan and Church, 1994; Haque et al., 2020a), or lead to misinterpretations (Chmutina et al., 2021; Yue et al., 2024).
@@
In the field of AI, terms such as “Coreference Resolution” or “Explain-Away Effect,” are frequently translated incorrectly or inconsistently, undermining comprehension for global researchers and practitioners (Khuwaileh and Khwaileh, 2011; Tehseen et al., 2018), as illustrated in Figure 1. For non-English readers, who comprise a substantial portion of the global population and AI community (Ammon, 2003; Ding et al., 2023), such inaccuracies in AI research paper and blog translation hinder access to essential knowledge, stifling research innovation and collaboration across linguistic boundaries (Amano et al., 2023; Bahji et al., 2023). The implications extend to AI resources documentation, such as model and data cards hosted on platforms like Hugging Face,111huggingface.co/docs/hub/en/model-cards, huggingface.co/docs/hub/en/datasets-cards where errors in translated terminologies can lead to misunderstandings or incorrect usage of models and datasets. This highlights the urgent need for precise and standardized multilingual AI terminology resources to support equitable access to AI knowledge (Ahuja et al., 2023; Liu et al., 2024).
@@
Existing efforts to address this gap have relied heavily on manual curation of terminologies by domain experts, a process that is time-consuming, resource-intensive, and difficult to scale (Wang et al., 2013; Freitag et al., 2021). Meanwhile, large language models (LLMs) have shown promise in automated terminology translation (Feng et al., 2024), while their outputs often misalign with human expert standards (Zhang et al., 2024), and different models often yield inconsistent translations (Banik et al., 2019; Prieto Ramos, 2021). The ACL 60-60 initiative in 2022 curated and translated AI-specific terms, and showcased potential for multilingual scientific communication and AI inclusivity (Salesky et al., 2023).2222022.aclweb.org/dispecialinitiative.html However, its small scale, narrow scope, and translation quality limitations underscore the need for a more comprehensive multilingual AI terminology resource.
@@
To enhance accessibility of our work, we integrate the curated terminologies into machine translation pipelines without requiring model retraining. We evaluate three post-translation refinement methods: prompting, word alignment and replacement, and constrained decoding. Our experiments reveal that the prompting method effectively integrates the curated dictionary, consistently improving translation quality as measured by BLEU and COMET scores. Additionally, we develop a website demonstration that incorporates our dataset into translations of ACL Anthology research papers, showcasing its practical applications for non-English speakers in the AI field.
@@
  • We created GIST, the first large-scale multilingual AI terminology dataset, with 5K terms from 879 awarded top AI conference papers (2000–2023) and a broad domain coverage. [^3] Our data and code have been uploaded to the submission system, and will be open-sourced upon paper acceptance.
  • We developed an effective and efficient translation framework combining LLMs and human expertise to translate English terms into five languages, achieving high-quality results validated by automatic and human evaluations.
  • To enhance accessibility, we explored three approaches to integrate the curated terminologies into machine translation pipelines without retraining, and developed a website to showcase its practical use for AI paper translation.
@@ 2 Related Works
Scientific terminology datasets are typically built using two approaches. The first approach relies on human multilingual experts for extracting and translating terminologies (Awadh, 2024; Kim et al., 2024). In cases of limited expert availability, human crowdsourcing with aggregation techniques has shown excellent results (Zaidan and Callison-Burch, 2011; Chan et al., 2023). The second approach uses ML and machine translation tools for automatic collection and translation, including log-likelihood comparisons (Haque et al., 2018), machine translation-based data synthesis (Haque et al., 2020b; Fernando et al., 2020; Manzini et al., 2022; Moslem et al., 2023), platform-based terminology linkers (Arcan et al., 2014), classifier training (Jin et al., 2013; Schumann and Martínez Alonso, 2018), and prompting LLMs (Nishio et al., 2024; Shamsabadi et al., 2024).
@@
Integrating newly collected domain terminology into machine translation systems has led to a variety of research efforts. One common approach involves training methods, such as augmenting training data with input-output pairs that include terminologies for supervised fine-tuning (Dinu et al., 2019; Niehues, 2021), or modifying model architectures to enhance terminology awareness (Dinu et al., 2019; Conia et al., 2024). However, these training-based approaches are inefficient for adapting to new terminologies, limiting their real-world applicability.
@@ 3 GIST Dataset Construction
We constructed GIST, a dataset comprising around 5K English AI terminologies and their translations into Arabic, Chinese, French, Japanese, and Russian. The basic lexical statistics of the dataset are presented in Table 1. We detail our terminology collection process in Section 3.1 and the translation procedure in Section 3.2.
@@
Table 1: Lexical statistics of the dataset across languages. “En” denotes English, and “Tgt” denotes the target language. Terms are tokenized into words using nltk.word_tokenize for English, Arabic, French, and Russian, jieba for Chinese, and MeCab Owakati for Japanese. Statistics with standard deviations are presented in Table 3.
3.1 Terminology Collection
@@
Our dataset follows the 60-60 initiative that aims to collect scientific terminologies in the AI field. We source AI terminology from two primary channels: AI research papers published online and existing AI terminology dictionaries. For the research papers, our objective is to compile a substantial number of terms from high-quality AI papers spanning a long time frame. To identify representative papers, a natural approach would be to crawl the most cited papers using online search engines or platforms. However, no tools or APIs currently enable this. Instead, we focused on awarded papers announced on the websites of top AI conferences. Specifically, we collected all awarded papers, spanning awards such as Best Paper, Outstanding Paper, and other recognitions from the venues listed in Table 4, and covering the years 2000 to 2023. This approach ensures comprehensive coverage of recognized research work. In total, we collected 879 paper PDFs from arXiv and other online repositories and processed them into text files using SciPDF. [^4] https://github.com/titipata/scipdf_parser As we analyze later in Section 5, this collection strategy provides broad terminology and domain coverage of AI terminologies.
@@
As supported by previous research (Xu et al., 2024; Dagdelen et al., 2024; Shamsabadi et al., 2024), recent LLMs such as LLaMA 3 (Dubey et al., 2024) have demonstrated strong capabilities in scientific terminology extraction tasks. While the definition of “AI” lacks a clear and universally agreed boundary, LLMs are trained on vast datasets that reflect human knowledge, enabling them to classify terms as AI-related based on their contextual relevance. Leveraging this, we utilized LLaMA-3-70B-Instruct to extract AI terminologies from award-winning papers, providing specific instructions to guide the extraction process.
@@
To define AI-specific terminology, we imposed the following criteria: (1) the term must be a noun or noun phrase, (2) it should be specialized to AI, encompassing core concepts, methods, models, algorithms, or systems, and (3) it should have either no meaning or a distinct meaning outside the AI domain. We processed the text in sentence chunks of up to 64 words to stay within LLaMA’s optimal context-handling capabilities. Additionally, for each unique term, we recorded up to three different contexts during extraction to ensure sufficient contextual diversity.
@@
After extraction, we performed multiple quality assurance steps. We removed terms that appeared in only one paper to ensure representativeness, excluded abbreviations and terms starting with special characters, and filtered out non-noun phrases. Duplicates were eliminated, and GPT-4o (Hurst et al., 2024) was employed to further refine the list by filtering out non-AI terms based on the same criteria. Finally, three AI domain experts reviewed the terms to remove any remaining unqualified entries. To enhance the dataset, we also integrated terminologies from external sources, including the 60-60 initiative dataset, government websites, Wikipedia, and other online resources. Full details of the terminology collection process are provided in Section A.2.
@@ 3.2 Terminology Translation
To evaluate whether state-of-the-art LLMs can effectively perform terminology translation as a generation task, we initially experimented with two advanced LLMs and one API available at the time, including Claude 3 Sonnet (Anthropic, 2024), GPT-3.5-Turbo (OpenAI, 2023), and Google Translate API (Wu et al., 2016). The procedure and the prompt used for these translations is presented in Section A.3. We measured the agreement among the three methods using exact match. However, as shown in Table 5, the three-model agreement ratio was only around 15% for most languages except Chinese, and the two-model agreement ratio was only about 40%. These results reveal significant inconsistencies in the translations produced by these systems, and highlight the need for human inputs to ensure reliable translations.
Given these findings, we opted for human annotation to ensure translation accuracy. To achieve this, we utilized Amazon Mechanical Turk (MTurk) for crowdsourced annotations. A demonstration of the MTurk task is shown in Figure 6. We instructed participants to take on the task only if they specialized in AI and were fluent in both English and one of the target languages. Annotators were tasked with generating accurate translations for each AI terminology, with relevant contexts provided and the terminology highlighted in yellow in contexts. To maintain quality, we implemented a rigorous qualification process. Annotators were first tested on a toy set of 10 carefully selected AI terms, and only those who performed well were allowed to proceed with the full task. Additionally, we monitored submissions daily and filtered out participants who provided random or low-quality translations during the annotation process. For each term, we collected 10 annotations per target language. The total cost is around \$1000.
Finally, we used GPT-4o to select the best translation from the annotators’ submissions for each term, ensuring high-quality results for our final dataset. As analyzed in Section 4.1, leveraging GPT-4o was crucial for maintaining the quality of the translations.
@@ 4 Dataset Quality Assessments
To thoroughly evaluate the translation quality in GIST, we conducted two additional crowdsourced rating tasks: (1) In Section 4.1, we investigated whether using GPT-4o was necessary to select the best translation candidate; (2) In Section 4.2, we compared the quality of our dataset with the evaluation set from the 60-60 initiative.
@@ 4.1 Task 1: Is Our LLM + Human Framework Necessary?
To ensure the selection of the best candidate from the annotators’ generations, we explored two methods for candidate selection. The first method only uses GPT-4o to select the best translation candidate among 10 annotations for all terms. The second method relies on majority voting among the 10 annotators: a translation candidate is selected if it appears in more than 5 out of 10 annotations. In cases where no majority is reached, GPT-4o is prompted to select the best translation candidate.
To evaluate these approaches, we conducted a separate MTurk task, involving a different group of participants, to compare the two methods on a randomly sampled subset of 200 terms per language. Participants were asked to choose one of four options: (A) Both translations are good; (B) Method 1 translation is better; (C) Method 2 translation is better; (D) Both translations are bad. For each language, we collected 5 annotations per term. As shown in Table 6, GPT-4o’s candidate selection consistently outperformed majority voting across all five languages. These results highlight the necessity of GPT-4o in achieving high-quality translations for crowdsourced annotations.

### v3

@@ Abstract
The field of machine translation has achieved significant advancements, yet domain-specific terminology translation, particularly in AI, remains challenging. We introduce GIST, a large-scale multilingual AI terminology dataset containing 5K terms extracted from top AI conference papers spanning 2000 to 2023. The terms are translated into Arabic, Chinese, French, Japanese, and Russian using a hybrid framework that combines LLMs for extraction with human expertise for translation. The dataset’s quality is benchmarked against existing resources, demonstrating superior translation accuracy through crowdsourced evaluation. GIST is integrated into translation workflows using post-translation refinement methods that require no retraining, where LLM prompting consistently improves BLEU and COMET scores. A web demonstration on the ACL Anthology platform highlights its practical application, showcasing improved accessibility for non-English speakers. This work aims to address critical gaps in AI terminology resources and fosters global inclusivity and collaboration in AI research.
@@
A Large-Scale Multilingual Terminology Dataset (GIST)
@@ 1 Introduction
Figure 2: Overview of GIST creation. AI terminology are extracted from awarded papers using an LLM and then combined with existing terminology dictionaries. After data cleansing, translations into five languages are generated via crowdsourcing, and an LLM is used to select the best candidate translation. Dataset quality is evaluated against two baselines: majority-vote-selected translations and translations from the 60-60 evaluation set, demonstrating the superior quality of GIST.
@@
The field of machine translation has made significant progress, with state-of-the-art models excelling across diverse tasks (Brown et al., 1990; Wu et al., 2016; Goyal et al., 2022; Haddow et al., 2022) and demonstrating effectiveness in translating between high-resource and low-resource languages (Yao and Wan, 2020; Costa-jussà et al., 2022; Ranathunga et al., 2023). Despite these successes, translating domain-specific scientific texts remains a persistent challenge, particularly for terminology translation (Cabré, 2010; Shuttleworth, 2014; Naveen and Trojovskỳ, 2024). General-purpose translation systems often falter in accurately translating specialized terminology, leading to loss of critical details (Dagan and Church, 1994; Haque et al., 2020a), or worse yet possibly leading to misinterpretations (Chmutina et al., 2021; Yue et al., 2024).
@@
In the field of AI, terms such as “Coreference Resolution” or “Explain-Away Effect,” are frequently translated incorrectly or inconsistently, undermining comprehension for global researchers and practitioners (Khuwaileh and Khwaileh, 2011; Tehseen et al., 2018), as illustrated in Figure 1. For non-English readers, who comprise a substantial portion of the global population as well as the AI community (Ammon, 2003; Ding et al., 2023), such inaccuracies in AI research papers and blog translation hinder access to essential knowledge, stifling research innovation and collaboration across linguistic boundaries (Amano et al., 2023; Bahji et al., 2023). The implications extend to AI resource documentation, such as model and data cards hosted on platforms such as Hugging Face,111huggingface.co/docs/hub/en/model-cards, huggingface.co/docs/hub/en/datasets-cards where errors in translated terminology can lead to misunderstandings or incorrect usage of models and datasets. This highlights the urgent need for precise and standardized multilingual AI terminology resources to support equitable access to AI knowledge (Ahuja et al., 2023; Liu et al., 2024).
@@
Existing efforts to address this gap have relied heavily on manual curation of specialized terminology by domain experts, a process that is time-consuming, resource-intensive, and difficult to scale (Wang et al., 2013; Freitag et al., 2021). Meanwhile, large language models (LLMs) have shown promise in automated terminology translation (Feng et al., 2024), while their outputs often misalign with human expert standards (Zhang et al., 2024), and different models often yield inconsistent translations (Banik et al., 2019; Prieto Ramos, 2021). The ACL 60-60 initiative in 2022 curated and translated AI-specific terms, and showcased potential for multilingual scientific communication and AI inclusivity (Salesky et al., 2023).2222022.aclweb.org/dispecialinitiative.html However, its small scale and narrow scope underscore the need for a more comprehensive multilingual AI terminology resource at scale.
@@
To enhance accessibility of our work, we integrate the curated terminology into machine translation pipelines without requiring model retraining. We evaluate three post-translation refinement methods: prompting, word alignment and replacement, and constrained decoding. Our experiments reveal that the prompting method effectively integrates the curated dictionary, consistently improving translation quality as measured by BLEU (Papineni et al., 2002) and COMET scores (Rei et al., 2020). Additionally, we develop a website demonstration that incorporates our dataset into translations of ACL Anthology research papers, showcasing its practical application for non-English speakers in the AI field.
@@
  • We create GIST, the first large-scale multilingual AI terminology dataset, with 5K terms from 879 awarded top AI conference papers (2000–2023) and broad domain coverage;333 Our code and data is available at https://github.com/jiarui-liu/MultilingualAITerminology.
  • We develop an effective and efficient translation framework combining LLMs and human expertise to translate English terms into five languages, achieving high-quality results validated by automatic and human evaluations.
  • To enhance accessibility, we explore three approaches to integrate the curated terminology into machine translation pipelines without retraining, and develop a website to showcase its practical use for AI paper translation.
@@ 2 Related Work
Scientific terminology datasets are typically built using two approaches. The first relies on human multilingual experts for extracting and translating terminology (Awadh, 2024; Kim et al., 2024). In cases of limited expert availability, human crowdsourcing with aggregation techniques has shown excellent results (Zaidan and Callison-Burch, 2011; Chan et al., 2023). The second approach uses ML and machine translation tools for automatic collection and translation, including log-likelihood comparisons (Haque et al., 2018), machine translation-based data synthesis (Haque et al., 2020b; Fernando et al., 2020; Manzini et al., 2022; Moslem et al., 2023), platform-based terminology linkers (Arcan et al., 2014), classifier training (Jin et al., 2013; Schumann and Martínez Alonso, 2018), and prompting LLMs (Nishio et al., 2024; Shamsabadi et al., 2024).
@@
Integrating newly collected domain terminology into machine translation systems has led to a variety of research efforts. One common approach involves training methods, such as augmenting training data with input-output pairs that include terminology for supervised fine-tuning (Dinu et al., 2019; Niehues, 2021), or modifying model architectures to enhance terminology awareness (Dinu et al., 2019; Conia et al., 2024). However, these training-based approaches are inefficient for adapting to new terminology, limiting their real-world applicability.
@@ 3 GIST Dataset Construction
We construct GIST, a dataset comprising around 5K English AI terminology and their translations into Arabic, Chinese, French, Japanese, and Russian. The basic lexical statistics of the dataset is presented in Table 1.
@@
Table 1: Statistics of the dataset across languages. “En” denotes English, and “Tgt” denotes the target language. Terms are tokenized into words using nltk.word_tokenize for English, Arabic, French, and Russian, jieba for Chinese, and MeCab Owakati for Japanese. Statistics with standard deviations are presented in Table 3.
3.1 Terminology Curation
@@
Our dataset follows the ACL 60-60 initiative that aims to collect scientific terminology in the AI field. We source AI terminology from two primary channels: AI research papers published online and existing AI terminology dictionaries. For the research papers, our objective is to compile a substantial number of terms from high-quality AI papers spanning a long time frame. To identify representative papers, a natural approach would be to crawl the most cited papers using online search engines or platforms. However, no tools or APIs currently enable this. Instead, we focus on awarded papers announced on the websites of top AI conferences. Specifically, we collect all awarded papers, spanning awards such as Best Paper, Outstanding Paper, and other recognitions from the venues listed in Table 4, covering the years 2000 to 2023. This approach ensures comprehensive coverage of recognized research work. In total, we collect 879 paper PDFs from arXiv and other online repositories and process them into text files using SciPDF. [^4] https://github.com/titipata/scipdf_parser As we analyze later in Section 5, this collection strategy provides broad terminology and domain coverage of AI terminology.
@@
As supported by previous research (Xu et al., 2024; Dagdelen et al., 2024; Shamsabadi et al., 2024), recent LLMs such as LLaMA 3 (Dubey et al., 2024) have demonstrated strong capabilities in scientific terminology extraction tasks. While the definition of “AI” lacks a clear and universally agreed boundary, LLMs are trained on vast datasets that reflect human knowledge, enabling them to classify terms as AI-related based on their contextual relevance. Accordingly, we leverage LLaMA-3-70B-Instruct to extract AI terminology from award-winning papers, providing specific instructions to guide the extraction process.
@@
To define AI-specific terminology, we impose the following criteria: (1) the term must be a noun or noun phrase, (2) it should be specialized to AI, encompassing core concepts, methods, models, algorithms, or systems, and (3) it should have either no meaning or a distinct meaning outside the AI domain. We process the text in sentence chunks of up to 64 words to stay within LLaMA’s optimal context-handling capabilities. Additionally, for each unique term, we record up to three different contexts during extraction to ensure sufficient contextual diversity.
@@
After extraction, we perform multiple quality assurance steps. We remove terms that appear in only one paper to ensure representativeness. Moreover, we exclude abbreviations and terms starting with special characters, and filter out non-noun phrases. Duplicates are eliminated, and GPT-4o (Hurst et al., 2024) is employed to further refine the list by filtering out non-AI terms based on the same criteria. Finally, three AI domain experts review the terms to remove any remaining unqualified entries. To enhance the dataset, we also integrate terminology from external sources, including the 60-60 initiative dataset, government websites, Wikipedia, and other online resources. Full details of the terminology collection process are provided in Section A.2.
@@ 3.2 Terminology Translation
In selecting target languages for translation, we aim to encompass a range of morphological complexities and varying levels of resource availability, particularly in AI-related publications. First, Chinese and Japanese exhibit minimal morphological variation in nouns and noun phrases, relying primarily on word order and context rather than inflection. In contrast, Arabic, French, and Russian are morphologically complex, characterized by extensive inflectional systems. Second, French, Chinese, Japanese, and Russian were selected due to the presence of large native-speaking scientific communities (Ammon, 2012; Chahal et al., 2022), which may have historically contributed to the development of well-established AI terminology. In contrast, Arabic, despite being widely spoken, may lack the same depth of scientific vocabulary, particularly in rapidly evolving fields such as AI. This disparity underscores the potential for our approach to contribute to the standardization and development of cohesive terminology across different languages in AI publications.
To evaluate whether state-of-the-art LLMs can effectively perform terminology translation as a generation task, we initially experiment with two advanced LLMs and one API: Claude 3 Sonnet (Anthropic, 2024), GPT-3.5-Turbo (OpenAI, 2023), and Google Translate API (Wu et al., 2016). The procedure and the prompt used for these translations is presented in Section A.3. We measure the agreement among the three methods using exact match. However, as shown in Table 5, the three-model agreement ratio was only around 15% for most languages except Chinese, and the two-model agreement ratio was only about 40%. These results reveal significant inconsistencies in the AI terminology translations produced by these systems, and highlight the need for human input to ensure reliable AI terminology translations.
Given these findings, we opt for human annotation to ensure translation accuracy. To achieve this, we utilize Amazon Mechanical Turk (MTurk) for crowdsourced annotations. A demonstration of the MTurk task is shown in Figure 6. We instruct participants to take on the task only if they specialized in AI and are fluent in both English and one of the target languages. Annotators are tasked with generating accurate translations for each AI terminology, with relevant contexts provided and the terminology highlighted in yellow in context. To maintain quality, we implement a rigorous qualification process. Annotators are first tested on a toy set of 10 carefully selected AI terms, and only those who perform well are allowed to proceed with the full task. Additionally, we monitor submissions daily and filtered out participants who provide random or low-quality translations during the annotation process. All in all, for each term, we collect 10 annotations per target language. The total cost is around \$1000.
Finally, we use GPT-4o to select the best translation from the annotators’ submissions for each term, ensuring high-quality results for our final dataset. As analyzed in Section 4.1, leveraging GPT-4o is crucial for maintaining the quality of the translations.
@@ 4 Dataset Quality Assessments
To thoroughly evaluate the translation quality in GIST, we conduct two additional crowdsourced rating tasks: (1) In Section 4.1, we investigate whether using GPT-4o is necessary to select the best translation candidate; (2) In Section 4.2, we compare the quality of our dataset with the evaluation set from the 60-60 initiative.
@@ 4.1 Task 1: Is Our LLM + Human Framework Necessary?
To ensure the selection of the best candidate from the annotators’ generations, we explore two methods for candidate selection. The first method only uses GPT-4o to select the best translation candidate among 10 annotations for all terms. The second method relies on majority voting among the 10 annotators: a translation candidate is selected if it appears in more than 5 out of 10 annotations. In cases where no majority is reached, GPT-4o is prompted to select the best translation candidate.
To evaluate these approaches, we conduct a separate MTurk task, involving a different group of participants, to compare the two methods on a randomly sampled subset of 200 terms per language. Participants are asked to choose one of four options: (A) Both translations are good; (B) Method 1 translation is better; (C) Method 2 translation is better; (D) Both translations are bad. For each language, we collect 5 annotations per term. As shown in Table 6, GPT-4o’s candidate selection consistently outperforms majority voting across all five languages. These results highlight the necessity of GPT-4o in achieving high-quality translations for crowdsourced annotations.